# Enhancing Group Fairness in Federated Learning through Personalization

### v1

@@ 4 Personalization can also improve fairness: numerical illustration
We begin by numerically illustrating the (unintended) fairness benefits of personalization, and elaborate on the reasons behind it. (We substantiate these with analytical support in Section 5.) We will compare the average local statistical parity (SP) fairness achieved by two personalized FL algorithms, Clustered-FL (Ghosh et al., 2020) and MAML-FL (Fallah et al., 2020), against FedAvg and standalone learning (where each client learns by itself, and there is no collaboration). Similar experiments supporting our findings under other notions of fairness (EqOp, EO) are given in Appendix E.
@@
We present numerical experiments on the “Retiring Adult” dataset (Ding et al., 2021). The dataset consists of census data collected from 50 U.S. states, and Puerto Rico. In our context, each state is a client within the FL framework. Each data sample includes multi-dimensional features $x$ (e.g., age, education, citizenship), a true label $y$, and a protected attribute $g$ (race, sex). To provide a more clear comparison of the effects of personalization, we have manually scaled the feature set ($x$) by 60% for the states with IDs {1, 10, 20, 30, 40, 50}; this exacerbates the data heterogeneity. We focus on two binary classification tasks: Employment (ACSEmployment) and Income (ACSIncome), and employ a two-layer neural network for both tasks. Each client is provided with 1000 training samples and 2000 testing samples. The values reported are the averages from 5 runs. Further details regarding the dataset and models, and additional figures, can be found in Appendix C.
@@ 4.1 Imbalanced groups: statistical advantages of collaboration
Fig 1(c) shows that all collaborative training algorithms (FedAvg, MAML-FL, Clustered-FL) achieve better local fairness (smaller gap) compared to standalone learning. This is due to the statistical benefits of collaboration: each client has limited samples in the non-White group, leading to poorly trained models with high local fairness gap (and low accuracy). In contrast, collaborative training in essence has access to more data, improving both metrics. For the same reason, the Clustered-FL algorithm, which partitions clients into two clusters, has (slightly) worse local fairness compared to FedAvg. In comparison, the MAML-FL algorithm, which effectively sees the global dataset (when training the global model that is later fine-tuned by each client), has better local fairness compared to FedAvg, indicating that personalization can improve both local accuracy (as intended) and local fairness (as a side benefit).
@@ 4.2 Better-balanced groups: computational advantages of collaboration
We first notice from Fig 2(c) that all collaborative training algorithms still have better local fairness compared to standalone learning. Furthermore, we observe that Clustered-FL and MAML-FL improve both local accuracy and local fairness compared to FedAvg. This is due to the computational advantages of (personalized) collaborative learning: for each client, due to similarity of the data for the male and female groups (as seen in Figure 2(b)) the objective of maximizing local accuracy is aligned with reducing the local fairness gap. Therefore, collaboration improves local fairness, with personalization further enhancing the model’s local accuracy and therefore its fairness.
We also note that (local) accuracy and fairness may not necessarily be aligned. Our next experiment shows that personalization can still improve fairness in such tasks compared to non-personalized FedAvg, which (we posit) is driven by a combination of statistical and computational benefits. Specifically, we conduct experiments on another task, ACSIncome, with “sex” as the protected attribute. Fig 3(a) shows that for this task, the fraction of samples is comparable across groups for label 0 data, but differs for label 1 data. From Fig 3(c), we observe that this time, all collaborative training algorithms improve accuracy but have worse local fairness compared to standalone learning; this is because improving (local) accuracy is not aligned with fairness in this task. That said, we observe that the personalized FL algorithms slightly improve local fairness compared to FedAvg. We interpret this as the statistical advantage of (effectively) observing more label 1 data (as FedAvg does, too), combined with a computational advantage of not overfitting a global model to the majority label 0 data (unlike what FedAvg may be doing).
@@ 4.3 Comparison with a fair Federated Learning algorithm
Figure 4: Comparing FedAvg, FairFed, Clustered-FL, MAML-FL, and standalone learning.
@@
To show this, we consider the FairFed algorithm of Ezzeldin et al. (2023), which adjusts the global model aggregation weights according to a fairness metric with the goal of improving (global) fairness. We also consider an extension of the algorithm, FairFed-AF, which adjusts the weights in aggregation according to both accuracy and fairness metrics. The experiments in Fig. 4 are on the ACSEmployment task with sex as the protected attribute. We observe that fair FL algorithms can achieve the best fairness across all algorithms. However, they have worse local accuracy compared to other collaborative training algorithms, as they focus (at least partially) on improving fairness and not accuracy.
5 Personalization can also improve fairness: theoretical support
To support and validate our findings from the numerical experiments in Section 4, in this section, we analytically show that personalized Federated clustering algorithms (which cluster/group similar clients to improve their models’ local accuracy) can also lead to better local fairness, when compared to a (non-personalized) shared global model.
Recall from the setup of Section 3 that we assume the $n$ clients can be potentially grouped into two clusters, $C_{\alpha}$ and $C_{\beta}$, based on their data similarities in their data distributions $f^{y,c}_{g}(x)$, with a fraction $p$ of clients in cluster $C_{\alpha}$. For personalized learning, we consider a cluster-based FL algorithm where each cluster can learn its own optimal cluster-specific model $\theta^{*}_{i},~{}i\in\{\alpha,\beta\}$ (obtained after solving equation 1), and contrast that with the average optimal model $\theta^{*}_{G}$ that would be obtained if all $n$ clients collaboratively learn a shared global model. We then contrast the average local fairness $\Delta^{\alpha}_{f}(\theta)$ obtained for clients in cluster $C_{\alpha}$ under a personalized model $\theta^{*}_{\alpha}$ vs. a shared model $\theta^{*}_{G}$, for two notions of fairness: $f\in\{\texttt{EqOp},\texttt{SP}\}$.
We start with the EqOp (Equality of Opportunity) fairness constraint, which aims to equalize true positive rates (TPR) between the protected groups $a$ and $b$. The following proposition shows that if $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$ (i.e., the data heterogeneity is such that cluster $C_{\alpha}$ has a lower optimal threshold than $C_{\beta}$), then clients in cluster $C_{\alpha}$ can obtain better local fairness (in addition to better local accuracy) with their cluster-specific model compared to if they used a global model shared with clients in $C_{\beta}$.
Proposition 1 (Improved EqOp through clustering).
Assume $f^{y,c}_{g}(x),y\in\{0,1\},g\in\{a,b\},c\in\{C_{\alpha},C_{\beta}\}$, are unimodal distributions, with modes $m^{y,c}_{g}$ such that $m^{y,c}_{b}\leq m^{y,c}_{a},\forall i,c$, and $\alpha^{1,c}_{g}\geq\alpha^{0,c}_{g},\forall g,c$. If $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$, there exist a cluster size $\hat{p}$ such that for $p\geq\hat{p}$, we have $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{\alpha})<\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{G})$; that is, the global model is more unfair than the cluster-specific model for $C_{\alpha}$.
The proof is presented in Appendix D.1. Intuitively, clients in $C_{\alpha}$ are better off under their personalized model as, given the proposition’s conditions, an increase in the decision threshold (which happens when moving from $\theta^{*}_{i}$ to $\theta^{*}_{G}$) will decrease the TPR of the disadvantaged group $b$ (the one with a lower mode in its feature distribution) slower than that of the advantaged group $a$, increasing the fairness gap for clients in $C_{\alpha}$.
@@ 5 Personalization can also improve fairness: theoretical support
We next consider SP (Statistical Parity) fairness, which assesses to disparity in the selection (positive classification) rate between the two protected groups. This is impacted by both the group $a$ vs. $b$ feature distributions as well as the label rates, rendering it more stringent than EqOp fairness. Figure 5 illustrates this by plotting the fairness gap vs. the decision threshold $\theta$ for SP vs. EqOp, showing that SP exhibits less structured changes as the decision threshold moves (e.g., due to the use of a global model). Therefore, to facilitate theoretical analysis, we further assume that $f^{y,c}_{g}(x)$ follows Gaussian distributions with equal variance $\sigma^{2}$ but different means $\mu^{y,c}_{g}$, such that $\mu^{0,c}_{b}\leq\mu^{0,c}_{a}\leq\mu^{1,c}_{b}\leq\mu^{1,c}_{a}$; the ordering is chosen so that label 0 samples have lower features than label 1 samples, and that for the same label, group $a$ samples have higher features than group $b$ samples (making group $a$ advantaged). We again find that then clients in cluster $C_{\alpha}$ can obtain better local SP-fairness (in addition to better local accuracy) with their cluster-specific model compared to if they joined in on a global model shared with clients in $C_{\beta}$.
Proposition 2 (Improved SP through clustering).
Assume $f^{y,c}_{g}(x),y\in\{0,1\},g\in\{a,b\},c\in\{C_{\alpha},C_{\beta}\}$, are Gaussian distributions with means $\mu^{0,c}_{b}\leq\mu^{0,c}_{a}\leq\mu^{1,c}_{b}\leq\mu^{1,c}_{a}$ and equal variance $\sigma^{2}$. Assume further that $\mu^{1,c}_{a}-\mu^{0,c}_{a}=\mu^{1,c}_{b}-\mu^{0,c}_{b}$, $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$, and that either $\alpha^{1,c}_{g}\geq\alpha^{0,c}_{g},\forall g$, or $r^{c}_{b}\geq r^{c}_{a},\forall c$ and $\alpha^{1,c}_{a}>\alpha^{0,c}_{b}>\alpha^{1,c}_{b}>\alpha^{0,c}_{a}$. Then, if $\alpha^{0,c}_{a}\exp(\frac{(\bar{\theta}-\mu^{0,c}_{a})^{2}}{-2\sigma^{2}})(\bar{\theta}-\mu^{0,c}_{a})-\alpha^{1,c}_{b}\exp(\frac{(\bar{\theta}-\mu^{1,c}_{b})^{2}}{-2\sigma^{2}})(\bar{\theta}-\mu^{1,c}_{b})\geq\alpha^{0,c}_{b}\exp(\frac{(\bar{\theta}-\mu^{0,c}_{b})^{2}}{-2\sigma^{2}})(\bar{\theta}-\mu^{0,c}_{b})-\alpha^{1,c}_{a}\exp(\frac{(\bar{\theta}-\mu^{1,c}_{a})^{2}}{-2\sigma^{2}})(\bar{\theta}-\mu^{1,c}_{a})$ holds, where $\bar{\theta}:=\tfrac{\mu^{1,c}_{a}+\mu^{0,c}_{b}}{2}$, there exist a $\hat{p}$ such that for $p\geq\hat{p}$, $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{\alpha})<\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{G})$.
A detailed proof is given in Appendix D.2. Proposition 2 assumes an equal distance between mean estimates (reflecting a uniform, systematic underestimation of group $b$ features); we relax this in Appendix E.7-E.10. Intuitively, the proposition states the following: when $\alpha^{1,c}_{g}>\alpha^{0,c}_{g}$, there are more label 1 data in both groups, and $\theta^{*}_{G}$ will pull $\theta^{*}_{\alpha}$ up to account for the label imbalance, resulting in a deterioration in both fairness and accuracy. Similarly, the other condition means that group $a$’s clients are majority label 1, while group $b$’s clients are majority label 0; then, $\theta^{*}_{\alpha}<\bar{\theta}$ if $r_{b}\geq r_{a}$, resulting in a higher fairness gap for clients in $\mathcal{C}_{\alpha}$ under $\theta^{*}_{G}$ for the same reason as the first condition.
6 Fair-FCA: A Fairness-aware Federated Clustering Algorithm
@@
We have observed numerically (Section 4) and analytically (Section 5) that integrating personalization techniques (such as clustering) in the FL paradigm can provide a dual benefit by not only enhancing local accuracy as intended, but also improving (local) group fairness. Motivated by this, we propose a new Fairness-aware Federated Clustering Algorithm (Fair-FCA), which builds on existing clustered FL ideas to further integrate fairness considerations in cluster identity assignment.
6.1 The Fair-FCA algorithm
@@
Our algorithm is an extension of the Clustered-FL (IFCA) algorithm of Ghosh et al. (2020), by modifying the cluster identity assignment step to encompass a fairness criterion. Specifically, we consider a FL scenario involving a total of $n$ clients and $K$ disjoint clusters. Each client $i$ has $n_{i}$ samples $Z_{i}:=\{z_{ij}\}^{n_{i}}_{j=1}$. Let $f(z,\theta)$ be the loss function associated with data point $z$ under model $\theta$. Then, for client $i$, the empirical loss is given by $F(Z_{i},\theta):=\frac{1}{n_{i}}\sum_{j}f(z_{ij},\theta)$. Additionally, let $\Psi^{f}(Z_{i},\theta)$ represent the fairness under model $\theta$ for fairness metric $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$ assessed on a dataset $Z_{i}$ (i.e., $\Delta_{f}(\theta)$ empirically calculated on $Z_{i}$).
@@
Our algorithm iterates over two steps: (1) cluster identity assignment, and (2) training of cluster-specific models. Specifically, let $\Theta^{t}_{k}$ denote cluster $k$’s model at time step $t$. The cluster identity for client $i$ at time $t$, denoted $c^{t}(i)$, is determined by:

### v2

@@ 4 Personalization can also improve fairness
We begin by numerically illustrating the (unintended) fairness benefits of personalization, and elaborate on the reasons behind it. (We substantiate these with analytical support in Section LABEL:sec:analysis.) We will compare the average local statistical parity (SP) fairness achieved by personalized FL algorithms, IFCA (ghosh2020efficient), FL+HC (briggs2020federated), CFL-Cosine (sattler2020clustered), MAML-FL (fallah2020personalized), and FedProx (li2020federated), against FedAvg (mcmahan2017communication) and Standalone learning. Similar experiments supporting our findings under other notions of fairness (EqOp, EO) are given in Appendix LABEL:app:numerical_other_notion.
@@
We first present numerical experiments on the “Retiring Adult” dataset (ding2021retiring). The dataset consists of census data collected from 50 U.S. states, and Puerto Rico. In our context, each state is a client within the FL framework. Each data sample includes multi-dimensional features $x$ (e.g., age, education, citizenship), a true label $y$, and a protected attribute $g$ (race, sex). To provide a more clear comparison of the effects of personalization, we have manually scaled the feature set ($x$) by 60% for the states with IDs {1, 10, 20, 30, 40, 50}; this exacerbates the data heterogeneity. We focus on two binary classification tasks: Employment (ACSEmployment) and Income (ACSIncome), and employ a two-layer neural network for both tasks. Each client is provided with 1000 training samples and 2000 testing samples. The values reported are the averages from 5 runs. Further details regarding the dataset and models, and additional figures, can be found in Appendix LABEL:app:_data_model. We also conduct similar numerical experiments under ACSHealth, another task in the Retiring Adult dataset, and find the same insights; these can be found in Appendix LABEL:app:numerical_other_data. We will also present numerical experiments on the “Adult” dataset (Dua:2019) in Section 4.3, and on synthetic data in Appendix LABEL:app:numerical-synthetic.
@@ 4.1 Imbalanced groups: statistical advantages of collaboration
Fig 1(c) shows that all collaborative training algorithms (FedAvg, MAML-FL, IFCA, FedProx, FL+HC, and CFL-Cosine) achieve better local fairness (smaller gap) compared to Standalone learning. This is due to the statistical benefits of collaboration: each client has limited samples in the non-White group, leading to poorly trained models with high local fairness gap (and low accuracy). In contrast, collaborative training in essence has access to more data, improving both metrics. For the same reason, the IFCA, FL+HC and CFL-Cosine algorithms, which partition clients into multiple clusters, has (slightly) worse local fairness compared to FedAvg. Similarly, the FedProx algorithm imposes a regularization term that prevents the local updates from deviating too much from the global model, making it less fair compared to FedAvg. In comparison, the MAML-FL algorithm, which effectively sees the global dataset (when training the global model that is later fine-tuned by each client), has better local fairness compared to FedAvg, indicating that personalization can improve both local accuracy (as intended) and local fairness (as a side benefit).
@@ 4.2 Better-balanced groups: computational advantages of collaboration
We first notice from Fig 2(c) that all collaborative training algorithms still have better local fairness compared to Standalone learning. Furthermore, we observe that all personalized learning algorithms (IFCA, FL+HC, CFL-Cosine, MAML-FL, and FedProx) improve both local accuracy and local fairness compared to FedAvg. This is due to the computational advantages of (personalized) collaborative learning: for each client, due to similarity of the data for the male and female groups (as seen in Figure 2(b)) the objective of maximizing local accuracy is aligned with reducing the local fairness gap. Therefore, collaboration improves local fairness, with personalization further enhancing the model’s local accuracy and therefore its fairness.
We also note that (local) accuracy and fairness may not necessarily be aligned. Our next experiment shows that personalization can still improve fairness in such tasks compared to non-personalized FedAvg, which (we interpret) is driven by a combination of statistical and computational benefits. Specifically, we conduct experiments on another task, ACSIncome, with “sex” as the protected attribute. Fig 3(a) shows that for this task, the fraction of samples is comparable across groups for label 0 data, but differs for label 1 data. From Fig 3(c), we observe that this time, all collaborative training algorithms improve accuracy but have worse local fairness compared to Standalone learning; this is because improving (local) accuracy is not aligned with fairness in this task. That said, we observe that the personalized FL algorithms slightly improve local fairness compared to FedAvg. We interpret this as the statistical advantage of (effectively) observing more label 1 data (as FedAvg does, too), combined with a computational advantage of not overfitting a global model to the majority label 0 data (unlike what FedAvg may be doing).
@@
4.3 Experiments on the Adult dataset
The Adult dataset (Dua:2019) is a widely used dataset from the UCI Machine Learning Repository, containing demographic data from the 1994 U.S. Census. Its task is to predict whether an individual’s income exceeds \$50k annually, based on a combination of features on the individual, including their demographics. The dataset consists of 48,842 samples described by 14 attributes, both categorical and numerical. Among these, 41,762 samples belong to the White group, while 7,080 samples are from the Non-White groups. Given the nature of this data’s heterogeneity, we randomly generate 5 clients, each with an unbalanced number of samples based on race. Additionally, in Appendix LABEL:app:numerical_other_data, we conduct experiments where samples are distributed with less heterogeneity across 5 clients, as done in other existing FL studies (e.g. (ezzeldin2023fairfed)).
(a) Number of samples
(b) Local accuracy vs. fairness gap
Figure 4: Experiments on the Adult dataset with race as the protected attribute.
From Fig 4(b), we can see that the results are consistent with our findings in Section 4.1. Interestingly, we observe that the IFCA algorithm clusters clients by grouping those with more White samples into one cluster and those with more Non-White samples into another (i.e., {0,1,2}, {3,4}). In contrast, the FL+HC clusters clients by grouping those with more samples into one cluster and those with less samples into another (i.e., {0,1,3,4}, {2}). As a result, these two variants of clustering-based algorithms have different performance, having statistical advantages but for different reasons.
@@ 4.4 Comparison with a fair Federated Learning algorithm
Figure 5: Comparing FedAvg, FairFed, IFCA, MAML-FL, FedProx, FL+HC, CFL-Cosine and Standalone learning.
@@
To show this, we consider the FairFed algorithm of ezzeldin2023fairfed, which adjusts the global model aggregation weights according to a fairness metric with the goal of improving (global) fairness. We also consider an extension of the algorithm, FairFed-AF, which adjusts the weights in aggregation according to both accuracy and fairness metrics. The experiments in Fig. 5 are on the ACSEmployment task with sex as the protected attribute. We observe that fair FL algorithms can achieve the best fairness across all algorithms. However, they have worse local accuracy compared to other collaborative training algorithms, as they focus (at least partially) on improving fairness and not accuracy.
4.5 Analytical support
To support and validate our findings from the numerical experiments, we also analytically show that personalized Federated clustering algorithms (which cluster/group similar clients to improve their models’ local accuracy) can also lead to better local fairness, when compared to a (non-personalized) shared global model, under certain conditions. Our results are detailed in Appendix LABEL:sec:analysis. Specifically, we consider a setting where data has single-dimensional, normally distributed features, leading to optimal threshold-based classifiers, and assume that clients can be put into two clusters $C_{\alpha}$ and $C_{\beta}$ based on similarities in their local datasets $\mathcal{D}_{i}$.
We start with the EqOp (Equality of Opportunity) fairness constraint, which aims to equalize true positive rates (TPR) between the protected groups $a$ and $b$. Our first result (Proposition LABEL:prop3) shows that if $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$ (i.e., the data heterogeneity is such that cluster $C_{\alpha}$ has a lower optimal threshold than $C_{\beta}$), then clients in cluster $C_{\alpha}$ can obtain better local fairness (in addition to better local accuracy) with their cluster-specific model compared to if they used a global model shared with clients in $C_{\beta}$.
@@ 4.5 Analytical support
Our second result (Proposition LABEL:prop1) considers the SP (statistical parity) constraint, which assesses the disparity in the selection (positive classification) rate between the two protected groups. This is impacted by both the group $a$ vs. $b$ feature distributions as well as the label rates, rendering it more stringent than EqOp fairness. Figure 6 illustrates this by plotting the fairness gap vs. the decision threshold $\theta$ for SP vs. EqOp, showing that SP exhibits less structured changes as the decision threshold moves (e.g., due to the use of a global model). Therefore, to facilitate theoretical analysis, we impose additional assumptions on the data distributions. Intuitively, the resulting proposition states the following: when there are more label 1 samples in both groups, the global model $\theta^{*}_{G}$ will pull the $C_{\alpha}$ cluster model $\theta^{*}_{\alpha}$ up to account for the label imbalance, resulting in a deterioration in both fairness and accuracy for clients in this cluster. Similarly, if group $a$ ($b$)’s clients are majority label 1 (0), then, the use of global instead of a local models can result in a higher fairness gap for clients in $\mathcal{C}_{\alpha}$ for the same reason as earlier.
5 Fairness-aware Federated Clustering Algorithms
@@
We have observed numerically (Section 4) and analytically (Appendix LABEL:sec:analysis) that integrating personalization techniques (such as clustering) in the FL paradigm can provide a dual benefit by not only enhancing local accuracy as intended, but also improving (local) group fairness. Motivated by this, we propose to further enhance the fairness of the learned personalized models by modify the clustering assignment process to also include fairness considerations.
We illustrate this idea by proposing new fairness-aware Federated clustering algorithms, Fair-FCA and Fair-FL+HC, based on the existing IFCA algorithm (ghosh2020efficient) and FL+HC algorithm (briggs2020federated), respectively. We choose these algorithms as many existing algorithms in the clustered FL literature are built on the IFCA framework (e.g., li2021federated; chung2022federated; huang2023active; ma2024structured) and the FL+HC framework (e.g., jothimurugesan2023federated; luo2023privacy; li2023hierarchical; sun2024collaborate.
5.1 Integrating fairness metrics in cluster identity assignment
@@
Consider a FL scenario involving a total of $n$ clients and $K$ disjoint clusters. Each client $i$ has $n_{i}$ samples $Z_{i}:=\{z_{ij}\}^{n_{i}}_{j=1}$. Let $f(z,\theta)$ be the loss function associated with data point $z$ under model $\theta$. Then, the empirical loss of client $i$ is given by $F(Z_{i},\theta):=\frac{1}{n_{i}}\sum_{j}f(z_{ij},\theta)$. Also, let $\Psi^{f}(Z_{i},\theta)$ denote the fairness under model $\theta$ for fairness metric $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$ assessed on $Z_{i}$.
The Fair-FCA algorithm.
@@
This algorithm iterates over two steps: (1) cluster identity assignment, and (2) training of cluster-specific models. Specifically, let $\Theta^{t}_{k}$ denote cluster $k$’s model at time step $t$. The cluster identity for client $i$ at time $t$, denoted $c^{t}(i)$, is determined by:
@@
The Fair-FL+HC algorithm.
Initially, the algorithm runs the regular FedAvg procedure for a predetermined number of rounds before clustering. Once a global model $\theta^{FA}$ is obtained, each client receives the model and performs several local updates to personalize their local models $\theta_{i}$.
Like the FL+HC algorithm, the Fair-FL+HC also employs hierarchical clustering with a set of hyperparameters $P$ to group clients by minimizing intra-cluster variance, measured using the $L_{2}$ Euclidean distance metric. We extend this approach by also considering fairness performance.
[…]
Here, $\mathbf{D}$ is a symmetric matrix where each entry $D_{i,j}$ represents the Euclidean distance between $\theta_{i}$ and $\theta_{j}$. $\mathbf{\Psi}^{f}(Z,\theta):=max(\Psi^{f}(Z_{i},\theta_{j}),\Psi^{f}(Z_{j},\theta_{i}))$, with $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$ is also a symmetric matrix that captures the worst-case $f$-fairness performance when client $i$’s model is evaluated on client $j$’s local data, or vice versa. The parameter $\gamma$ balances between fairness and accuracy considerations; when $\gamma=1$, we recover the FL+HC algorithm. Once clustering is completed, each cluster trains its model independently using FedAvg. The pseudo-code for Fair-FL+HC is shown in Algorithm 2.
5.2 Fairness-accuracy tradeoff using Fair-FCA and Fair-FL+HC
Figure 7: Acc. vs. SP-fairness
We begin by conducting a numerical experiment on a synthetic dataset, to illustrate the ability of Fair-FCA and Fair-FL+HC to strike a balance between fairness and accuracy. We consider a total of 8 clients that could (potentially) be clustered into two clusters. Among these, 6 clients (Client ID: 2,4,5,6,7,8) have similar data distributions, with 4 clients (Client ID: 4,6,7,8) sharing identical distributions across the two protected groups $a,b$ (low fairness gap). The remaining 2 clients (Client ID: 1,3) have different data distributions compared to the first six, but they also share identical distributions across the two protected groups. We consider $f=\texttt{SP}$. Let $\gamma_{1},\gamma_{2}\in[0,1]$ be the hyperparameters of the Fair-FCA and Fair-FL+HC algorithms, respectively. Additionally, let the distance threshold $d\in P$ in Fair-FL+HC be 0.35.
Figure 7 shows that when $\gamma_{1}=\gamma_{2}=1$, both Fair-FCA and Fair-FL+HC prioritizes accuracy; by design, this is attained by grouping the 6 clients having similar data distributions together ({1,3} and {2,4,5,6,7,8}). Similarly, when $\gamma_{1}=\gamma_{2}=0$, both Fair-FCA and Fair-FL+HC focus only on SP fairness, this time clustering clients that have identical distributions on the two protected groups together ({2,5} and {1,3,4,6,7,8}). Lastly, by setting $\gamma_{1},\gamma_{2}\in(0,1)$, we can effectively account for both accuracy and SP fairness when clustering: when $\gamma_{1}=0.3,\gamma_{2}=0.1$, the clusters are {2,4,5} and {1,3,6,7,8}; when $\gamma_{1}=0.5,\gamma_{2}=0.3$, the clusters are {2,4,5,6} and {1,3,7,8}; and when $\gamma_{1}=0.8,,\gamma_{2}=0.6$, the clusters are {2,4,5,6,7} and {1,3,8}.
5.3 Comparing Fair-FCA vs. IFCA; Fair-FL+HC vs. FL+HC
We now compare the IFCA and FL+HC algorithms of ghosh2020efficient and briggs2020federated, respectively, and our Fair-FCA and Fair-FL+HC algorithms with $\gamma=0.5$ and with $f=\texttt{SP}$, in terms of local accuracy and local group fairness. We proceed with the same experiment setting as in Section 4. (We also provide experiment results using the original Retiring Adult dataset without feature scaling in Appendix LABEL:app_experiment_original_retiring, and experiments with EqOp as the fairness metric in Appendix LABEL:app_experiment_eqop_retiring.)
Figure 8 verifies that under different datasets (ACSEmployment, ACSIncome) and different protected attributes (race, sex), Fair-FCA and Fair-FL+HC (symbols in squares and triangles) exhibits a fairness-accuracy tradeoff (improved fairness performance at the expense of accuracy degradation) consistent with the choice of $\gamma=0.5$ compared to the IFCA and FL+HC algorithm.
Moreover, we can observe that FL+HC could achieve better accuracy compared to IFCA in the ACSEmployment tasks with both race and sex attributes. This is because IFCA uses a predetermined number of clusters, set to 2 in all experiments. As a result, it effectively groups scaled clients into one cluster and unscaled ones into the other. In contrast, FL+HC employs a hierarchical clustering technique, which does not require prior knowledge of the number of clusters. This allows FL+HC not only to successfully group scaled clients into one cluster but also to form additional clusters (if needed) for other different clients. The fairness performance of IFCA and FL+HC remains consistent with our findings in Section 4.
(a) Employment (Race)
(b) Employment (Sex)
(c) Income (Sex)
Figure 8: Fair-FCA and Fair-FL+HC with different tasks and protected attributes ($f=\texttt{SP}$)